The osmotic pressure of an aqueous solution at 330. K is 115 kPa. Estimate the freezing point of the solution. (The freezing-point constant of H2O is 1.86 K·kg·mol−1. Assume the mass density of H2O is 1000. kg·m−3.)
1
Expert's answer
2019-11-18T06:27:55-0500

P = C*R*T, where P = 115 000 Pa - the osmotic pressure R = 8,314 J*K-1*mol-1 - universal gas constant, T =330 K - temperature, C - molar concentration, mol/m3. From this equation: C = P/(R*T).= 115 000/(8,314*330) = 41,92 mol/m3.

The freezing point of the solution can be calculated by the formula: T = TH2O - ΔT, where TH2O = 0 oC.

ΔT = C*k/"\rho" = 41,92*1,86/1000 = 0,078 oC.

T = 0 - 0,078 = -0,078 oC